Mickey is working on building a piston. He hes worried that it cant withstand the 3) needed pressure because it seems to fail at 2.6atm of pr non-compressed stage at 600K it has a volume of 9L and a pressure of 1.6atm. When it essure. Will it work if in the compresses he removes heat to 300K and compresses it to 3L?
Answer #1

3) T = 600 K

V = 9 L

P = 1.6 atm

Vf = 3 L

Tf = 300 K

hence

from ideal gas equation

PV/T = PfVf/Tf

1.6*9/600 = Pf*3/300

Pf = 2.4 atm

as the piston fails at 2.6 atm, hence th epsiton will work in this case